OpenStudy (anonymous):

Still having problems with this one. 8(z-3)=2(z+12) Is this the right answer? So do I add the 8z and 2z together?

OpenStudy (anonymous):

What was the original question?

OpenStudy (anonymous):

8(z-3)=2(z+12)

OpenStudy (anonymous):

z = 8

OpenStudy (anonymous):

Need to know once I *8z and 2z do I do I add the 8 and 2?

OpenStudy (anonymous):

8(z-3)=2(z+12) 8Z-24=2Z+24 8Z-2Z =24+24 6Z=48 Z=8

OpenStudy (anonymous):

8*5=2*20

OpenStudy (anonymous):

If you're trying to solve for z you have to distribute the 8 and the 2. 8(z-3)=2(z+12) => 8z-24=2z+24 Then solve for z: subtract 2z from each side and add 24 to each side 6z=48 Divide each side by 6: z=8

OpenStudy (anonymous):

Thanks I get it now

OpenStudy (anonymous):

\[8(z-3)=2(z+12) \]distributive property \[8z-24=2z+24\] subtract 2z from both sides \[8z-2z-24=2z-2z+24\] \[6z-24=24\] add 24 to both sides \[6z-24+24=24+24\] \[6z=48\] divide both sides by 6 \[\frac{6z}{6}=\frac{48}{6}\] \[z=8\]

myininaya (myininaya):

\[8(z-3)=2(z+12)\] I would divide 2 on both sides first \[4(z-3)=z+12\] Now I would distribute 4(z-3)=4z-12 \[4z-12=z+12\] Subtract z on both sides \[3z-12=12\] Add 12 on both sides \[3z=24\] Divide 3 on both sides \[z=8\]